hhanan:

What is the standard form for an exponential equation? Explain how you know if an exponential equation shows growth or decay.

CripQUEZZ:

Wait what are these questions for

CripQUEZZ:

Like a test a question hw or what?

hhanan:

my assignment

CripQUEZZ:

Oh ok I jus didn’t wanna be helping on a test bc that’s not allowed

CripQUEZZ:

Dang I can’t figure it out

mxddi3:

hmm, maybe ask @Convert for the standard form, bc i don't remember. I have it in my notes that it's written as \[a \times (b)^x\] but i'm not 100% sure. For the how you can tell if it's a growth or decay--If the b is greater than one, we have a growth and if it's less than one, we have a decay. Let me show you some examples. `Growth:` \[f(x)=1500(1.074)^x\] our "b"--what's in the parentheses--is greater than one. `Decay:` \[f(x)=200(0.5)^x\] 0.5<1 so it's a decay. hopefully this helps, i'd just ask Convert about the standard form.

CripQUEZZ:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @mxddi3 hmm, maybe ask @Convert for the standard form, bc i don't remember. I have it in my notes that it's written as \[a \times (b)^x\] but i'm not 100% sure. For the how you can tell if it's a growth or decay--If the b is greater than one, we have a growth and if it's less than one, we have a decay. Let me show you some examples. `Growth:` \[f(x)=1500(1.074)^x\] our "b"--what's in the parentheses--is greater than one. `Decay:` \[f(x)=200(0.5)^x\] 0.5<1 so it's a decay. hopefully this helps, i'd just ask Convert about the standard form. \(\color{#0cbb34}{\text{End of Quote}}\) I think that’s a pretty good explanation

Convert:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @mxddi3 hmm, maybe ask @Convert for the standard form, bc i don't remember. I have it in my notes that it's written as \[a \times (b)^x\] but i'm not 100% sure. For the how you can tell if it's a growth or decay--If the b is greater than one, we have a growth and if it's less than one, we have a decay. Let me show you some examples. `Growth:` \[f(x)=1500(1.074)^x\] our "b"--what's in the parentheses--is greater than one. `Decay:` \[f(x)=200(0.5)^x\] 0.5<1 so it's a decay. hopefully this helps, i'd just ask Convert about the standard form. \(\color{#0cbb34}{\text{End of Quote}}\) Yeah what you said looks good :). The equation would be f(x) = a(b)^x, where a is the initial amount, and b is the growth factor or amount grown but cannot be 1 because 1^1 is always 1.